# Pseudo-Stereo Inputs: A Solution to the Occlusion Challenge in Self-Supervised Stereo Matching

### v1

@@ V Conclusion
In this work, we propose a self-supervised stereo matching paradigm with the core idea of a pseudo-stereo inputs strategy. This method effectively addresses the occlusion issue, which has consistently hindered the performance of direct self-supervised stereo matching. Both qualitative and quantitative experiments demonstrate a significant improvement over existing direct self-supervised methods based on photometric consistency.
In future research, we aim to extend this methodology to other related challenges. For instance, similar occlusion issues are also present in self-supervised optical flow estimation problems. The notable effectiveness of decoupling input and feedback images in stereo matching suggests a potential solution for addressing occlusions in self-supervised optical flow estimation as well. In addition to this, large visual models have now become crucial for further enhancing the capabilities of learning-based methods. Their reliance on vast amounts of training data has made the effective utilization of easily accessible unannotated data an urgent requirement. The self-supervised method proposed in this paper offers new insights into exploiting information from unlabeled stereo image pairs, which may be referenced for further application in these areas.

### v2

@@ 5 Conclusion
In this work, we propose a self-supervised stereo matching paradigm with the core idea of a pseudo-stereo inputs strategy. This method effectively addresses the occlusion issue, which has consistently hindered the performance of self-supervised stereo matching. Extensive experiments validate our method’s success in handling occlusions and demonstrate its robustness and generality across diverse datasets and network backbones.
Despite our significant progress on occlusions, this work does not propose explicit solutions for other challenges, such as textureless and reflective regions. In fact, unlike occlusions, providing accurate feedback for these ill-posed regions is inherently intractable with binocular cues alone. A promising solution is the multi-frame self-supervised paradigm, as explored in [33]. This creates a powerful synergy: temporal consistency across frames is key to resolving ambiguities in areas like textureless regions, while the accuracy of the single-frame disparity estimation is core to resolving depth ambiguities of moving objects. The substantially enhanced reliability of disparity estimation achieved by our work provides a much stronger foundation for this paradigm, creating significant potential to better leverage multi-frame information and tackle the remaining challenges in self-supervised 3D perception.